What must be the distance between point charge q1 26.0 mC and point charge q2 47.0 mC for the electrostatic force between them to have a magnitude of 5.70 N?

Grade:12th

2 Answers

Arun
6 years ago
Dear Abdul
 

We can use Coulomb’s law to figure out the distance between the two point charges. Coulomb’s law states:

F=k\dfrac{(q_1)(q_2)}{r^2}F=kr2(q1)(q2)

 

(Where FF is the electrostatic force, kk is Coulomb’s constant, q_1q1 is the charge of the first particle, q_2q2 is the charge of the second particle, and rr is the distance between the two charges)

 

Substitute the values we know into our equation.

F=k\dfrac{(q_1)(q_2)}{r^2}F=kr2(q1)(q2)

5.7=(8.99\times 10^9)\dfrac{(26\times 10^{-6})(47\times 10^{-6})}{r^2}5.7=(8.99×109)r2(26×106)(47×106)

 

(isolate for r)

5.7r^2=(8.99\times 10^9)(26\times 10^{-6})(47\times 10^{-6})5.7r2=(8.99×109)(26×106)(47×106)

r^2=1.9273298r2=1.9273298

r=\sqrt{1.9273298}=1.39r=1.9273298=1.39 m

Rohit
6 years ago
Use Coulomb`s law F= k×(q×q)/r^2Substitute the values u will get the answer.Note q and q are different charges they may have same or different magnitude.Regards.